OpenStudy (anonymous):

Which value(s) of x should be excluded for the polynomial? x^2-16/x^2-6x+9 A. 3 B. 3 and -3 C. 4 D. 4 and -4

OpenStudy (anonymous):

I think it's D

OpenStudy (istim):

Why excluded?

OpenStudy (anonymous):

wat do u mean?

OpenStudy (istim):

Actually, just ignore me. Someone else, help!

OpenStudy (anonymous):

you can never divide by 0 so where will there be a problem with this one?

OpenStudy (anonymous):

it's A- (3)

OpenStudy (anonymous):

because???

OpenStudy (anonymous):

ik how did u get tht?

OpenStudy (anonymous):

he is wrong, by the way

OpenStudy (anonymous):

okay so then wat is it philip

OpenStudy (anonymous):

you cannot divide by zero, so what values for x will cause problems here? that is the question

OpenStudy (anonymous):

what is the denominator? and when is it 0?

OpenStudy (anonymous):

sorry, forgot to ask, do you want to learn this stuff or just want to fill in some homework? I am ok with it either way.

OpenStudy (anonymous):

look Factorize them (x^2-16)/(x^2-6x+9) so (x^2-16) = (x-4)(x+4) (x^2-6x+9)=(X-3)(x-3) therefore (x-4)(x+4)/(x-3)(x-3) since any denominator makes the whole equation undefined Therefore the no. 3 is Refused i mean Answer (A)

OpenStudy (anonymous):

fill in sum homework but kind of want to know a little about it.

OpenStudy (anonymous):

i mean when the denominator=0 the whole equation is undefined

OpenStudy (anonymous):

yep, that's right. I was mistaken. appologies where appropriate

OpenStudy (anonymous):

okay thnk u jop and philip u guys helped aa lot